OpenStudy (anonymous):

Simplify the expression. x^2/3 y^-1/4 over x^1/2 y^-1/2

hero (hero):

\[\frac{x^{2/3}y^{-1/4}}{x^{1/2}y^{-1/2}}\]

OpenStudy (anonymous):

Yes, that. I tried posting that but apparently it didn't go through. Can you explain how to do it?

OpenStudy (anonymous):

@Hero

hero (hero):

Hang on

hero (hero):

Basically, in general negative exponents do not mean negative number. They mean "inverse" as in "fractional reciprocal". In other words something like: \(y^{-x}\) "means the the multiplicative inverse of \(y^x\)" or simply \(\dfrac{1}{y^x}\) So in summary, \(y^{-x} = \dfrac{1}{y^x}\)

hero (hero):

Similarly \(\dfrac{1}{y^{-x}} = y^x\)

hero (hero):

In other words, if you see a negative fractional exponent in the numerator, you can move the expression to the denominator and remove the negative and it will be an equivalent form.

hero (hero):

Thus: \(\dfrac{x^{2/3}y^{-1/4}}{x^{1/2}y^{-1/2}} = \dfrac{x^{2/3}y^{1/2}}{x^{1/2}y^{1/4}}\)

OpenStudy (anonymous):

I just reverse the y?

hero (hero):

At this point, what you want to do is use this rule of exponents: If b > c, then \(\dfrac{a^b}{a^c} = a^{b - c}\) or If b < c, then \(\dfrac{a^b}{a^c} = \dfrac{1}{a^{c - b}}\)

hero (hero):

The reason for this is because you want to avoid negative exponents at all costs during simplification. So always subtract the smaller exponent from the bigger exponent.

hero (hero):

In this case, 2/3 > 1/2 and 1/2 > 1/4 therefore, by rule of exponents, we can simplify the expression to: \(\dfrac{x^{2/3}y^{-1/4}}{x^{1/2}y^{-1/2}} = \dfrac{x^{2/3}y^{1/2}}{x^{1/2}y^{1/4}} = x^{(2/3 - 1/2)}y^{(1/2 - 1/4)}\) Which, at this point, simplifying becomes a matter of simply subtracting fractions.

OpenStudy (anonymous):

And that would be \[x ^{1/6}y ^{1/4}\]

OpenStudy (anonymous):

correct?